Tim and Judy mix two kinds of feed for pedigreed dogs. They wish to make 35 pounds of feed worth $0.36 per pound by mixing one kind worth $0.18 per pound with another worth $0.53 per pound. How many pounds of the cheaper kind should they use in the mix?

Answer by mananth(16949)

$ ---------------- quantity
Type I 0.18 ---------------- x pounds
Type II 0.53 ------ 35 - x pounds
Mixture 0.36 ---------------- 35

0.18x+0.53(35-x)=0.36*35
0.18x+0.53(35-x )=12.6
0.18x+ 18.55-0.53x=12.6
0.18x-0.53x=12.6--18.55
-0.35 x = -5.95
/ -0.35
x=17 pounds$ 0.18 Type I
18 pounds$ 0.53 Type II

m.ananth@hotmail.ca